# HOT: An Efficient Halpern Accelerating Algorithm for Optimal Transport Problems

### v1

@@ Abstract
This paper proposes an efficient HOT algorithm for solving the optimal transport (OT) problems with finite supports. We particularly focus on an efficient implementation of the HOT algorithm for the case where the supports are in $\mathbb{R}^{2}$ with ground distances calculated by $L_{2}^{2}$-norm. Specifically, we design a Halpern accelerating algorithm to solve the equivalent reduced model of the discrete OT problem. Moreover, we derive a novel procedure to solve the involved linear systems in the HOT algorithm in linear time complexity. Consequently, we can obtain an $\varepsilon$-approximate solution to the optimal transport problem with $M$ supports in $O(M^{1.5}/\varepsilon)$ flops, which significantly improves the best-known computational complexity. We further propose an efficient procedure to recover an optimal transport plan for the original OT problem based on a solution to the reduced model, thereby overcoming the limitations of the reduced OT model in applications that require the transport map. We implement the HOT algorithm in PyTorch and extensive numerical results show the superior performance of the HOT algorithm compared to existing state-of-the-art algorithms for solving the OT problems.
@@
Figure 1: Selected examples of color transfer based on the reduced optimal transport model with the optimal transport map recovered by Algorithm 1.
@@ I Introduction
ThE Kantorovich-Wasserstein (KW) distance has become a primer choice for measuring the similarity between two probability distributions. It has demonstrated remarkable success in various applications, including color transfer [1, 2, 3], texture synthesis and mixing [4], registration and warping [5], transport-based morphometry [6], and hypothesis testing [7], among others. Despite its powerful geometric framework for comparing probabilities, the KW distance is computationally expensive in general [8, 9]. Specifically, it requires solving an optimal transport (OT) problem, which is a (large-scale) linear programming (LP) in a discrete setting. Standard methods, such as the simplex method and the interior point method, suffer from high computational complexity relative to the problem size. Furthermore, these methods are difficult to parallelize, which can hardly benefit from the modern powerful graphics processing units (GPUs). Consequently, solving the LP problem of the discrete OT problem remains daunting in modern data-driven applications due to high computational and memory costs. This paper addresses these two challenges by proposing an efficient and easily parallelizable algorithm for solving an equivalent reduced model of the OT problem.
@@ I-A Related work and existing challenges
When it comes to computing the KW distance between two discrete probability distributions with $M$ supports, there are mainly two popular approaches: (i) Computing an approximated KW distance by solving the optimal transport problem with an additional entropy regularization [10]; (ii) Computing the KW distance via solving the corresponding LP problem [11]. The readers can refer to [9] and the references therein for a more detailed discussion of the algorithms for solving OT problems. Before introducing our new algorithm, we briefly discuss the challenges of the aforementioned approaches.
@@
Challenges with the Entropy-regularized approach: Due to its scalability, the Sinkhorn algorithm, and its improved versions have been widely adopted to compute an approximation of the KW distance in applications [10, 12, 13, 14]. In particular, the Sinkhorn algorithm can efficiently solve the regularized OT problem when the regularization parameter is moderate (i.e., no less than $10^{-2}$). However, a high-quality approximation of the KW distance is important for better performance in many applications, which requires solving the regularized OT problem with a small regularization parameter. Unfortunately, a small regularization parameter will usually cause numerical issues and a slower convergence for the Sinkhorn algorithm. Some stabilized and rescaling techniques [15] have been proposed to improve the robustness of solving the regularized OT problems, but the efficiency of the stabilized algorithms is unsatisfactory compared to the Sinkhorn algorithm.
@@
Challenges with the LP approach: Along this line, the interior point method [16] and the network simplex method [17, 18] are popular choices for obtaining solutions with high accuracy to the moderate scale LP problem. Recently, a semismooth Newton based inexact proximal augmented Lagrangian method [19] has been proposed for solving linear programming problems which can solve the OT problem as a special case. The semismooth Newton based algorithm can exploit the sparsity of the solution by the generalized Jacobian and show superior numerical performance compared to Gurobi in some examples. However, these solvers are not applicable for solving the problem on a very large scale due to the high computational complexity. The urgent need to solve large-scale OT problems in applications inspires extensive research in designing first-order methods. Among them, the Douglas-Rachford splitting algorithm [20] and primal-dual hybrid gradient method (PDHG) [21, 22, 23, 24]. Jambulapati et al. [25] proposed an algorithm based on a dual-extrapolation algorithm to achieve an $\widetilde{O}({M}^{2}/\varepsilon)$ computational complexity bound for obtaining an $\varepsilon$-approximate solution (in terms of objective function value)  [^1] Despite its better complexity bound, the empirical performance of this algorithm is not as efficient as other algorithms, such as accelerated gradient method [13, 14].. Recently, Zhang et al. [26] proposed an efficient Halpern-Peaceman-Rachford (HPR) algorithm for solving the OT model and the Wasserstein barycenter problem, which can obtain an $\varepsilon$-approximate solution (in terms of the Karush-Kuhn-Tucker (KKT) residual) of the OT model in $O({M}^{2}/\varepsilon)$ flops. We summarize some known complexity results for solving the OT problem in Table I. Readers can refer to [27, 26, 28] and the references therein for a more detailed discussion.
@@
Algorithm Complexity result
Sinkhorn [13] $\widetilde{O}({M^{2}C^{2}}/{\varepsilon^{2}})$
APDAGD [13, 29] $\widetilde{O}(M^{2.5}C/\varepsilon)$
Greenkhorn [29] $\widetilde{O}(M^{2}C^{2}/\varepsilon^{2})$
Accelerated Sinkhorn [29] $\widetilde{O}(M^{7/3}C^{4/3}/\varepsilon^{4/3})$
AAM [14] $\widetilde{O}(M^{2.5}C/\varepsilon)$
Dual extrapolation [25] $\widetilde{O}(M^{2}C/\varepsilon)$
HPD [30] $\widetilde{O}({M^{2.5}C}/{\varepsilon})$
HPR [26] ${O(M^{2}R}/{\varepsilon})$
HOT (This paper) $\mathbf{O(M^{1.5}R/\varepsilon)}$
Beyond the challenges above in computational efficiency, all these algorithms for the original OT problem with ${M}$ supports require at least a memory cost of $O({M}^{2})$. This memory cost makes it forbidden to compute the KW distance of two distributions with massive supports (i.e., the OT problem for computing the KW distance of two $512\times 512$ grey images has more than $6.8\times 10^{10}$ variables). This motivates researchers to explore possible reduced models for the OT problem. When the ground distances between supports in $\mathbb{R}^{d}$ are calculated by $L_{1}$-norm, Ling and Okada proposed an equivalent reduced model with $O(dM)$ memory cost to calculate the earth mover distance (equivalent to the KW distance) [31]. Recently, Auricchio et al. [32] extended the idea to the case where the ground distances between supports are calculated by $L_{2}^{2}$-norm and proposed an equivalent reduced model with $O(dM^{\frac{d+1}{d}})$ memory cost. The authors in [32] adopted the Network Simplex method to solve the reduced model and demonstrated superior performance in terms of computational and memory efficiency compared to the Sinkhorn and the improved Sinkhon method [3, 33] on examples of moderate scale. Unfortunately, the efficiency becomes unsatisfactory for very large-scale problems (see Section IV for details). Moreover, the transport plan is not available if we solve the reduced model, which is critical for a wide class of applications, such as color transfer[1, 34] and texture synthesis [4].
@@ I-B Contributions
Motivated by the recent advancements in the accelerated algorithms based on Halpern iteration [35, 26, 36, 37, 38, 39], we propose an efficient Halpern accelerating method for solving the reduced OT problem, which is abbreviated as “HOT” for convenience, to address the challenges in computing the KW distance with a finite number of supports. We particularly focus on an efficient implementation of the HOT algorithm for the case where the supports are in $\mathbb{R}^{2}$ with ground distances calculated by $L_{2}^{2}$-norm, which includes a wide class of applications as aforementioned. Specifically, HOT adopts a first-order algorithm with Halpern acceleration to solve the equivalent reduced OT model, which can obtain an $\varepsilon$-approximate solution in $O(1/\varepsilon)$ iterations [35, 26]. More importantly, we design a fast procedure for solving the subproblems with linear time complexity by fully exploiting the problem structure. This also makes the popular alternating direction method of multiplier (ADMM) [40, 41] scalable for solving the reduced OT model. Overall, our proposed HOT algorithm can compute an $\varepsilon$-approximation of the KW distance between two histograms supported on $M=m\times n$ bins within $O((m^{2}n+n^{2}m)/\varepsilon)$ flops. This is the best-known computational complexity for computing an approximate KW distance to our knowledge. Moreover, we propose an efficient algorithm to recover a transport plan based on the obtained solution of the reduced model, which releases the power of the reduced model in applications. We implement HOT in PyTorch and extensive numerical results will be shown in Section IV to demonstrate the superior and robust performance of HOT for computing the KW distance, compared to state-of-the-art algorithms, including Sinkhorn [10], Improved Sinkhorn [3, 33], Network Simplex method [17, 18], ADMM [40, 41], interior point method (in Gurobi).
@@ II-A An equivalent reduced model of the OT problem
With these notations and assumptions, the discrete OT problem for computing KW distance between histograms $\mu^{1}$ and $\mu^{2}$ can be defined as follows:
$\begin{array}[]{ll}\min\limits_{\pi}&\sum\limits_{(i,j)\in\mathcal{I}}\sum\limits_{(k,l)\in\mathcal{I}}c_{i,j;k,l}\pi_{i,j;k,l}\\ \text{s.t.}&\begin{aligned} &\sum\limits_{(k,l)\in\mathcal{I}}\pi_{i,j;k,l}=\mu^{1}_{i,j},\quad\forall(i,j)\in\mathcal{I},\\ &\sum\limits_{(i,j)\in\mathcal{I}}\pi_{i,j;k,l}=\mu^{2}_{k,l},\quad\forall(k,l)\in\mathcal{I},\\ &\pi_{i,j;k,l}\geq 0,\quad\forall(i,j)\in\mathcal{I},\ \forall(k,l)\in\mathcal{I},\end{aligned}\end{array}$ (1)
@@
In this paper, we focus on the case where $p=2$. By exploiting the separable structure of the ground distance, Auricchio et al. [32] proposed the following equivalent model in terms of the optimal objective function value:
$\begin{array}[]{ll}\min\limits_{f^{(1)},f^{(2)}}&\sum\limits_{(i,j)\in\mathcal{I}}\left[\sum\limits_{k=1}^{m}(k-i)^{2}f_{i,k,j}^{(1)}+\sum\limits_{l=1}^{n}(j-l)^{2}f_{k,j,l}^{(2)}\right]\\ \text{s.t.}&\begin{aligned} &\sum_{i=1}^{m}f_{i,k,j}^{(1)}=\sum_{l=1}^{n}f_{k,j,l}^{(2)},&\forall(k,j)\in\mathcal{I},\\ &\sum_{k=1}^{m}f_{i,k,j}^{(1)}=\mu^{1}_{i,j},&\forall(i,j)\in\mathcal{I},\\ &\sum_{j=1}^{n}f_{k,j,l}^{(2)}=\mu^{2}_{k,l},&\forall(k,l)\in\mathcal{I},\\ &f_{i,k,j}^{(1)}\geq 0,f_{k,j,l}^{(2)}\geq 0,&\forall(i,j),(k,l)\in\mathcal{I},\end{aligned}\end{array}$ (3)
@@
Similar to [42, Lemma 7.1], we can obtain that $A$ defined in (6) has full row rank, and
@@ II-B Reconstruct the transport plan from the reduced model
The absence of the transport plan $\pi$ makes the reduced OT model less favorable in applications where the transport map is necessary (i.e., color transfer [1, 2, 3]). We address this issue by proposing a fast algorithm (shown in Algorithm 1) to reconstruct an optimal transport plan of the original model from an optimal solution of the reduced model (3).
@@
Output: An optimal transport mapping $\pi$ of problem (1).
@@
The following proposition shows that the output of Algorithm 1 is an optimal transport plan for of the original OT model.
@@
Additionally, since
$f^{(1)}_{i,k,j}\leq\sum_{i=1}^{m}f_{i,k,j}^{(1)}=\sum_{l=1}^{n}f_{k,j,l}^{(2)},$
by fixing $i,j,k$, we obtain from Algorithm 1 that
@@
According to [32, Theorem 1], we know that the optimal objective function values of problems (1) and (3) are equivalent. Therefore, $\pi$ is an optimal solution to problem (1). ∎

### v2

@@ Abstract
This paper proposes an efficient HOT algorithm for solving the optimal transport (OT) problems with finite supports. We particularly focus on an efficient implementation of the HOT algorithm for the case where the supports are in $\mathbb{R}^{2}$ with ground distances calculated by $L_{2}^{2}$-norm. Specifically, we design a Halpern accelerating algorithm to solve the equivalent reduced model of the discrete OT problem. Moreover, we derive a novel procedure to solve the involved linear systems in the HOT algorithm in linear time complexity. Consequently, we can obtain an $\varepsilon$-approximate solution to the optimal transport problem with $M$ supports in $O(M^{1.5}/\varepsilon)$ flops, which significantly improves the best-known computational complexity. We further propose an efficient procedure to recover an optimal transport plan for the original OT problem based on a solution to the reduced model, thereby overcoming the limitations of the reduced OT model in applications that require the transport plan. We implement the HOT algorithm in PyTorch and extensive numerical results show the superior performance of the HOT algorithm compared to existing state-of-the-art algorithms for solving the OT problems.
@@
Figure 1: Selected examples of color transfer based on the reduced optimal transport model with the optimal transport plan recovered by Algorithm 1.
@@ I Introduction
ThE Kantorovich-Wasserstein (KW) distance has become a prime choice for measuring the similarity between two probability distributions. It has demonstrated remarkable success in various applications, including color transfer [pitie2007linear, bonneel2013example, solomon2015convolutional], texture synthesis and mixing [dominitz2009texture], registration and warping [haker2004optimal], transport-based morphometry [basu2014detecting], and hypothesis testing [del1999tests], among others. Despite its powerful geometric framework for comparing probabilities, the KW distance is computationally expensive in general [villani2003topics, peyre2019computational]. Specifically, it requires solving an optimal transport (OT) problem, which is a (large-scale) linear programming (LP) in a discrete setting. Standard methods, such as the simplex method and the interior point method, suffer from high computational complexity relative to the problem size. Furthermore, these methods are difficult to parallelize, which can hardly benefit from the modern powerful graphics processing units (GPUs). Consequently, solving the LP problem of the discrete OT problem remains daunting in modern data-driven applications due to high computational and memory costs. This paper addresses these two challenges by proposing an efficient and easily parallelizable algorithm for solving an equivalent reduced model of the OT problem.
@@ I-A Related work and existing challenges
When it comes to computing the KW distance between two discrete probability distributions with $M$ supports, there are mainly two popular approaches: (i) Computing an approximated KW distance by solving the optimal transport problem with an additional entropy regularization [cuturi2013sinkhorn]; (ii) Computing the KW distance via solving the corresponding LP problem [orlin1988faster]. The readers can refer to [peyre2019computational] and the references therein for a more detailed discussion of the algorithms for solving OT problems. Before introducing our new algorithm, we briefly discuss the challenges of the aforementioned approaches.
@@
Challenges with the Entropy-regularized approach: Due to its scalability, the Sinkhorn algorithm and its improved versions have been widely adopted to compute an approximation of the KW distance in applications [cuturi2013sinkhorn, lin2019efficient, dvurechensky2018computational, guminov2021combination]. In particular, the Sinkhorn algorithm can efficiently solve the regularized OT problem when the regularization parameter is moderate (i.e., no less than $10^{-2}$). However, a high-quality approximation of the KW distance is important for better performance in many applications, which requires solving the regularized OT problem with a small regularization parameter. Unfortunately, a small regularization parameter will usually cause numerical issues and a slower convergence for the Sinkhorn algorithm. Some stabilized and rescaling techniques [schmitzer2019stabilized] have been proposed to improve the robustness of solving the regularized OT problems, but the efficiency of the stabilized algorithms is unsatisfactory compared to the Sinkhorn algorithm.
@@
Challenges with the LP approach: Along this line, the interior point method [pele2009fast] and the network simplex method [goldberg1989network, gabow1991faster] are popular choices for obtaining solutions with high accuracy to the moderate scale LP problem. Recently, a semismooth Newton based inexact proximal augmented Lagrangian method [li2020asymptotically] has been proposed for solving linear programming problems which can solve the OT problem as a special case. The semismooth Newton based algorithm can exploit the sparsity of the solution by the generalized Jacobian and show superior numerical performance compared to Gurobi in some examples. However, these solvers are not applicable for solving the problem on a very large scale due to the high computational complexity. The urgent need to solve large-scale OT problems in applications inspires extensive research in designing first-order methods, such as the Douglas-Rachford splitting algorithm [mai2022fast] and primal-dual hybrid gradient method (PDHG) [zhu2008efficient, esser2010general, chambolle2011first, applegate2021practical]. Jambulapati et al. [jambulapati2019direct] proposed an algorithm based on a dual-extrapolation algorithm to achieve an $\widetilde{O}({M}^{2}/\varepsilon)$ computational complexity bound, where $M$ is the number of supports, for obtaining an $\varepsilon$-approximate solution (in terms of objective function value) [^1] Despite its better complexity bound, the empirical performance of this algorithm is not as efficient as other algorithms, such as accelerated gradient method [dvurechensky2018computational, guminov2021combination].. Recently, Zhang et al. [zhang2022efficient] proposed an efficient Halpern-Peaceman-Rachford (HPR) algorithm for solving the OT model and the Wasserstein barycenter problem, which can obtain an $\varepsilon$-approximate solution (in terms of the Karush-Kuhn-Tucker (KKT) residual) of the OT model in $O({M}^{2}/\varepsilon)$ flops. We summarize some known complexity results for solving the OT problem in Table I. Readers can refer to [tupitsa2022numerical, zhang2022efficient, khamis2024scalable] and the references therein for a more detailed discussion.
@@
Algorithm Time complexity result Space complexity result
Sinkhorn [dvurechensky2018computational] $\widetilde{O}({M^{2}C^{2}}/{\varepsilon^{2}})$ ${O}({M^{2}})$
APDAGD [dvurechensky2018computational, lin2022efficiency] $\widetilde{O}(M^{2.5}C/\varepsilon)$ ${O}({M^{2}})$
Greenkhorn [lin2022efficiency] $\widetilde{O}(M^{2}C^{2}/\varepsilon^{2})$ ${O}({M^{2}})$
Accelerated Sinkhorn [lin2022efficiency] $\widetilde{O}(M^{7/3}C^{4/3}/\varepsilon^{4/3})$ ${O}({M^{2}})$
AAM [guminov2021combination] $\widetilde{O}(M^{2.5}C/\varepsilon)$ ${O}({M^{2}})$
Dual extrapolation [jambulapati2019direct] $\widetilde{O}(M^{2}C/\varepsilon)$ ${O}({M^{2}})$
HPD [chambolle2022accelerated] $\widetilde{O}({M^{2.5}C}/{\varepsilon})$ ${O}({M^{2}})$
HPR [zhang2022efficient] [^2] Zhang, G., Yuan, Y., & Sun, D. (2022). An Efficient HPR Algorithm for the Wasserstein Barycenter Problem with $O({\rm Dim(P)}/\varepsilon)$ Computational Complexity. arXiv preprint arXiv:2211.14881. (Under review at JMLR) ${O(M^{2}R}/{\varepsilon})$ ${O}({M^{2}})$
HOT (Ours) $\mathbf{O(M^{1.5}R/\varepsilon)}$ $\mathbf{O(M^{1.5})}$
Beyond the challenges above in computational efficiency, all these algorithms for the original OT problem with ${M}$ supports require at least a memory cost of $O({M}^{2})$. This memory cost makes it forbidden to compute the KW distance of two distributions with massive supports (i.e., the OT problem for computing the KW distance of two $512\times 512$ grey images has more than $6.8\times 10^{10}$ variables). To address this issue, researchers have proposed some approximate models for the OT problem, such as the linear OT framework [wang2013linear], the sliced Wasserstein distance [bonneel2011displacement], and the approximated earth mover’s distance [indyk2003fast, shirdhonkar2008approximate, leeb2016holder]. While these approaches alleviate memory issues and simplify computations, they inevitably introduce a loss of exactness. This limitation has motivated further research into equivalent reduced models for the OT problem. When the ground distances between supports in $\mathbb{R}^{d}$ are calculated by $L_{1}$-norm, Ling and Okada proposed an equivalent reduced model with $O(dM)$ memory cost to calculate the earth mover distance (equivalent to the KW distance) [ling2007efficient]. Recently, Auricchio et al. [auricchio2018computing] extended the idea to the case where the ground distances between supports are calculated by $L_{2}^{2}$-norm and proposed an equivalent reduced model with $O(dM^{\frac{d+1}{d}})$ memory cost. The authors in [auricchio2018computing] adopted the Network Simplex method to solve the reduced model and demonstrated superior performance in terms of computational and memory efficiency compared to the Sinkhorn and the convolutional Sinkhorn method [solomon2015convolutional, solomon2018optimal] on examples of moderate scale. Unfortunately, the efficiency becomes unsatisfactory for very large-scale problems (see Section IV for details). Moreover, the transport plan is not available if we solve the reduced model, which is critical for a wide class of applications, such as color transfer [pitie2007linear, ferradans2014regularized], texture synthesis [dominitz2009texture], and domain adaptation [courty2016optimal].
@@ I-B Contributions
Motivated by the recent advancements in the accelerated algorithms based on Halpern iteration [sun2024accelerating, zhang2022efficient, kim2021accelerated, tran2021halpern, lieder2021convergence, yang2025accelerated], we propose an efficient Halpern accelerating method for solving the reduced OT problem, which is abbreviated as “HOT” for convenience, to address the challenges in computing the KW distance with a finite number of supports. We particularly focus on an efficient implementation of the HOT algorithm for the case where the supports are in $\mathbb{R}^{2}$ with ground distances calculated by $L_{2}^{2}$-norm, which includes a wide class of applications as aforementioned. Specifically, HOT adopts a first-order algorithm with Halpern acceleration to solve the equivalent reduced OT model, which can obtain an $\varepsilon$-approximate solution in $O(1/\varepsilon)$ iterations [sun2024accelerating, zhang2022efficient]. More importantly, we design a fast procedure for solving the subproblems with linear time complexity by fully exploiting the problem structure. This also makes the popular alternating direction method of multipliers (ADMM) [glowinski1975approximation, gabay1976dual] scalable for solving the reduced OT model. Overall, our proposed HOT algorithm can compute an $\varepsilon$-approximation of the KW distance between two histograms supported on $M=m\times n$ bins within $O((m^{2}n+n^{2}m)/\varepsilon)$ flops. This is the best-known computational complexity for computing an approximate KW distance to our knowledge. Moreover, we propose an efficient algorithm to recover a transport plan based on the obtained solution of the reduced model, which releases the power of the reduced model in applications. We implement HOT in PyTorch and extensive numerical results will be shown in Section IV to demonstrate the superior and robust performance of HOT for computing the KW distance, compared to state-of-the-art algorithms, including Sinkhorn [cuturi2013sinkhorn], convolutional Sinkhorn[solomon2015convolutional, solomon2018optimal], Network Simplex method [goldberg1989network, gabow1991faster], ADMM [glowinski1975approximation, gabay1976dual], interior point method (in Gurobi).
@@ II-A An equivalent reduced model of the OT problem
With these notations and assumptions, the discrete OT problem for computing the KW distance between histograms $\mu^{1}$ and $\mu^{2}$ can be defined as follows:
$\begin{array}[]{ll}\min\limits_{\pi}&\sum\limits_{(i,j)\in\mathcal{I}}\sum\limits_{(k,l)\in\mathcal{I}}c_{i,j;k,l}\pi_{i,j;k,l}\\ \text{s.t.}&\left\{\begin{aligned} \quad&\sum\limits_{(k,l)\in\mathcal{I}}\pi_{i,j;k,l}=\mu^{1}_{i,j},\quad\forall(i,j)\in\mathcal{I},\\ &\sum\limits_{(i,j)\in\mathcal{I}}\pi_{i,j;k,l}=\mu^{2}_{k,l},\quad\forall(k,l)\in\mathcal{I},\\ &\pi_{i,j;k,l}\geq 0,\quad\forall(i,j)\in\mathcal{I},\ \forall(k,l)\in\mathcal{I},\end{aligned}\right.\end{array}$ (1)
@@
In this paper, we focus on the case where $p=2$. By exploiting the separable structure of the ground distance, Auricchio et al. [auricchio2018computing] proposed the following equivalent model in terms of the optimal objective function value:
$\begin{array}[]{ll}\min\limits_{f^{(1)},f^{(2)}}&\sum\limits_{(i,j)\in\mathcal{I}}\left[\sum\limits_{k=1}^{m}(k-i)^{2}f_{i,k,j}^{(1)}+\sum\limits_{l=1}^{n}(j-l)^{2}f_{k,j,l}^{(2)}\right]\\ \text{s.t.}&\left\{\begin{aligned} \quad&\sum_{i=1}^{m}f_{i,k,j}^{(1)}=\sum_{l=1}^{n}f_{k,j,l}^{(2)},&\forall(k,j)\in\mathcal{I},\\ &\sum_{k=1}^{m}f_{i,k,j}^{(1)}=\mu^{1}_{i,j},&\forall(i,j)\in\mathcal{I},\\ &\sum_{j=1}^{n}f_{k,j,l}^{(2)}=\mu^{2}_{k,l},&\forall(k,l)\in\mathcal{I},\\ &f_{i,k,j}^{(1)}\geq 0,f_{k,j,l}^{(2)}\geq 0,&\forall(i,j),(k,l)\in\mathcal{I},\end{aligned}\right.\end{array}$ (3)
@@
Similar to [dantzig2003linear, Lemma 7.1], we can obtain that $A$ defined in (6) has full row rank, and
@@ II-B Reconstruct the transport plan from the reduced model
The absence of the transport plan $\pi$ makes the reduced OT model less favorable in applications where the transport plan is necessary (i.e., color transfer [pitie2007linear, bonneel2013example, solomon2015convolutional]). We address this issue by proposing a fast algorithm (shown in Algorithm 1) to reconstruct an optimal transport plan of the original model from an optimal solution of the reduced model (3).
@@
Output: An optimal transport plan $\pi$ of problem (1).
@@
The following proposition shows that the output of Algorithm 1 is an optimal transport plan for the original OT model.
@@
Fix $i,j,k$. For convenience, let $f^{(1)-l}_{i,k,j}$ denote the value of $f^{(1)}_{i,k,j}$ at the $l$-th iteration before updating. According to the update formula of $f^{(1)}_{i,k,j}$, there must exist an index $1\leq l^{*}\leq n$ such that $f^{(1)-l^{*}}_{i,k,j}\leq f_{k,j,l^{*}}^{(2)}$. Otherwise, we would have
$f^{(1)}_{i,k,j}>\sum_{l=1}^{n}f^{(2)}_{k,j,l},$
which contradicts the feasibility of $(f^{(1)},f^{(2)})$:
$f^{(1)}_{i,k,j}\leq\sum_{i=1}^{m}f_{i,k,j}^{(1)}=\sum_{l=1}^{n}f_{k,j,l}^{(2)}.$
Therefore, after $n$ iterations, the final value of $f^{(1)}_{i,k,j}$ must be zero, which leads to
@@
According to [auricchio2018computing, Theorem 1], we know that the optimal objective function values of problems (1) and (3) are equivalent. Therefore, $\pi$ is an optimal solution to problem (1). ∎
Remark 1.
The worst-case computational complexity of reconstructing the transport plan via Algorithm 1 is $3M^{2}$. In practice, we can efficiently parallelize the $(k,j)$ loop to leverage the significant benefits of GPU acceleration, enabling an efficient reconstruction of the transport plan from a solution to the reduced OT model. Furthermore, in many applications, such as image retrieval [indyk2003fast] and shape matching [ling2007efficient], only the KW distance is required. In these scenarios, the reconstruction of the transport plan is not necessary.